OpenStudy (anonymous):

can this be simplified: -5.3xy + 9yx ?

OpenStudy (johnweldon1993):

yes it can...the order of the variables does not matter

OpenStudy (anonymous):

3.7xy

OpenStudy (johnweldon1993):

so since we have 2 numbers with the same variables they CAN be combined so -5.3xy + 9yx what is -5.3 + 9? 3.7 so 3.7xy

OpenStudy (anonymous):

3.7xy i think is this one but i never done one like that also gives me a choice that cannot be simplified

OpenStudy (anonymous):

the order of the variables in this case does not matter: xy = yx so you can do this: xy(9-5.3) = yx(9-5.3) = 3.7xy = 3.7yx

OpenStudy (anonymous):

alright thanks mate